OpenStudy (anonymous):

Can someone explain to me why 2/5-3=2/5-15/5?

OpenStudy (anonymous):

15/5=3 so when you subtract either of them from 2/5, you get the same answer.

OpenStudy (anonymous):

I don't understand why 3=15/5. I'm very confused as to why the whole number becomes that fraction and how do you know what the fraction value will be.

OpenStudy (amistre64):

15/5 means ... 15 divided by 5 does 15 divided by 5 = 3?

OpenStudy (anonymous):

\[\frac{15}{5}\Longleftrightarrow 15\div 5\]

OpenStudy (anonymous):

OHH OK!!! So the fraction will always be whatever the division problem would be to get the whole number?

OpenStudy (amistre64):

we also know that 3*1 = 3 and that 5/5 = 1 \[\frac{3}{1}*\frac{5}{5} = \frac{15}{5} = 3\]

OpenStudy (amistre64):

yes; fractions are just dividing

OpenStudy (amistre64):

12/3 = 4 18/9 = 2

OpenStudy (anonymous):

You are a God sent!!! I have been so stressed over this work!!! Would you mind working through one problem with me so I know what I am doing?

OpenStudy (amistre64):

post it and lets see

OpenStudy (amistre64):

and if you have any white out handy; just erase that "group master" jargon .... its soo misleading

OpenStudy (anonymous):

Lol. I was so confused by that!

OpenStudy (anonymous):

Find the equation of the line having the given slope and containing the given point. m=6 (9,1). The equation of the line in slope-intercept form is y=?

OpenStudy (anonymous):

y=6x-53

OpenStudy (anonymous):

it is just an lcm concept (2/5) + 3 on taking lcm it becomes (2 + 15)/5 ,then you split it into two 2/5 +15/5 hence they are equal

OpenStudy (anonymous):

its simple\[3*5\div1*5\]

OpenStudy (amistre64):

the point P(x,y), in this case (9,1) and the slope m = 6, here y = m(x-Px)+Py

OpenStudy (anonymous):

y=2x+8?

OpenStudy (amistre64):

y = 6(x-(9))+1 y = 6x -6(9) + 1 y = 6x -54 +1 y = 6x -53 is what I get for point intercept form

OpenStudy (anonymous):

Forget what I just typed, that was the example problem. I get mixed up having to go back and forth!

OpenStudy (anonymous):

y=6x-53?

OpenStudy (amistre64):

m= 6; P(9,1); slope intercept form: y = 6x -53 by the steps i took

OpenStudy (anonymous):

The next one is harder...

OpenStudy (amistre64):

lol ..... bring it on!! :)

OpenStudy (anonymous):

m=8/9, (8,-3)

OpenStudy (amistre64):

not harder; just different numbers :) y = mx -mPx +Py perhaps :)

OpenStudy (amistre64):

y = (8/9)x -(8/9)*8 -3

OpenStudy (anonymous):

would it be y+3=8/9x-4/2? Or am I wrong again on the fraction?

OpenStudy (amistre64):

y = (8/9)x -64/9 -27/9 y=(8/9)x - 91/9

OpenStudy (anonymous):

Why was it 91/9?

OpenStudy (amistre64):

this is the part your having difficulties with \[-(\frac{8}{9}*\frac{8}{1})-(\frac{3}{1}*\frac{9}{9})\]

OpenStudy (amistre64):

all ths amounts to is this: \[\frac{-64-27}{9}\]

OpenStudy (anonymous):

I feel like an idiot because my brain just won't wrap around the fraction part of it!

OpenStudy (amistre64):

its important to remember that in order to add or subtract fractions; they need to be wearing the same pants.... otherwise they dont want to talk to each other...

OpenStudy (amistre64):

multiplying fractions? they are fine with that .... but to join the group? you gotta have the right clothes on

OpenStudy (amistre64):

what number can we multiply by that doesnt change the value of a number?

OpenStudy (amistre64):

8 * ? = 8?

OpenStudy (anonymous):

1

OpenStudy (amistre64):

thats right; 1 seems pretty useless; but it really how we WRITE '1' that matters

OpenStudy (anonymous):

So figure out the simplified fraction?

OpenStudy (amistre64):

does anything over itself = 1? 3/3 = 1 ? 125/125 = 1 ? b/b = 1 ?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so we have a way in which we can "multiply" by '1' and change the way a fraction LOOKS; but keeps the same value right?

OpenStudy (anonymous):

Thank you dearly for your patience with me!

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so lets use that to our advantage :) - 8 8 -64 --- * -- = ---- right? 9 1 9

OpenStudy (anonymous):

Yes!

OpenStudy (amistre64):

but we get stuck here; becasue they arent wearing the same pants right? -64 3 ---- - --- 9 1

OpenStudy (anonymous):

This is where I get lost. How do you know that they aren't wearing the same pants? Again, sorry for all the questions. I just really need to learn and understand everything.

OpenStudy (amistre64):

-64/9 has on a 9 under it; its pants are a size '9' -64/9 ; pants are properly called a denominator 3/1 has a 1 under it; its pants are a size '1' right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

so lets see if there is any outfits in the closet to redress the 3/1 in to match the pants of -64/9 1 * ? = 9 ??

OpenStudy (anonymous):

9

OpenStudy (amistre64):

good; we have something that fits :) 3 9 27 -- * -- = --- <-- now we are dressed to mingle with the -64/9 1 9 9

OpenStudy (amistre64):

-64 -27 -64-27 -91 --- + ---- = ------- = ---- <-- this is true right? 9 9 9 9

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

do you see how that happened tho?

OpenStudy (anonymous):

so you can multiply something with the one that doesn;t fit to make to match?

OpenStudy (amistre64):

yes.... fractions will multiply like rabbits with no regard for what they look like.....

OpenStudy (anonymous):

Ok, do you mind moving onto the next problem or do you need to go?

OpenStudy (amistre64):

as long as you are getting the principals underneath this fraction stuff; im good to go to the next one :)

OpenStudy (anonymous):

OK, here we go.... :)

OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-2,-4) (-8,-1)

OpenStudy (amistre64):

'an' equation? no particular form?

OpenStudy (anonymous):

I got a slope of -3/6. Is that right?

OpenStudy (anonymous):

m=y2-y1/x2-x1= is the formula

OpenStudy (amistre64):

slope = (change in y)/(change in x) and I like to stack my points to keep track of them: ( x , y ) (-2,-4) -(-8,-1) -------- ( 6, -3); slope = y/x = -3/6 = -1/2

OpenStudy (anonymous):

So always simplify the answer?

OpenStudy (amistre64):

always? ..... most math tests and class WANT you to so yeah.

OpenStudy (amistre64):

-3/3 -1 ----- = ---- 6/3 2

OpenStudy (anonymous):

So then we would move onto, y-(-1)=-1/2(x-(-8))

OpenStudy (amistre64):

yes; for point slope form of a line that would be good

OpenStudy (anonymous):

y+1=-1/2x+4?

OpenStudy (amistre64):

.... -4

OpenStudy (anonymous):

-1/2x8/1=8/2 or 4?

OpenStudy (amistre64):

-1 8 -8 --- * --- = ---- = -4 2 1 2

OpenStudy (anonymous):

y=-1/2x-3?

OpenStudy (amistre64):

close.... y+1 = (-1/2)x -4 -1 -1 ----------------- y =(-1/2)x -5

OpenStudy (anonymous):

y=-1/2x-5!

OpenStudy (amistre64):

thats better :)

OpenStudy (anonymous):

Write an equation in standard form of the line passing through the given pair of points. The line goes through (2,4) (2,2). what is the equation? x+y=4 x=2 y=4 or undefined

OpenStudy (anonymous):

A whole new equation to confuse me, lol.

OpenStudy (amistre64):

well; I can see that the x parts are the same; so it really doesnt matter what the y parts are.... (2,y) satisfies this; so, x=2

OpenStudy (anonymous):

So if the x's are the same, it would just be x=whatever the number is?

OpenStudy (amistre64):

thats right; since x always = that number; the equation is: x = that number :)

OpenStudy (anonymous):

:) I am beginning to get a little more confident!

OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (9,-2); 6x-7y=4. the equation of the line is y=?

OpenStudy (anonymous):

Would I plug in the points of the x and y in the equation 6x-7y=4 or 6(9)-7(-2)=4?

OpenStudy (amistre64):

parralel points hav ethe same equation; just a different 'sum'; so ignore the "=4" part and fill in your x and y values to get a new one

OpenStudy (anonymous):

Wow! I am wrong on this one!

OpenStudy (amistre64):

6(9) -7(-2) = new sum for parallel line 54 +14 = n 68 = n 6x -7y = 68

OpenStudy (amistre64):

you probably missed your sign

OpenStudy (anonymous):

Now, I have to figure out the slope so I can continue to y=?

OpenStudy (amistre64):

why ? what form of the line equation you need?

OpenStudy (anonymous):

The help topics first suggestive step is 6x-7y=4 should be to subtract 6x from both sides.

OpenStudy (anonymous):

So -7y=4-6x?

OpenStudy (amistre64):

6x -7y = 68 is an equation for a line y+2 = -6/7(x-9) is the same line y = -6/7x +68/7 is the same line 6x 7y --- - --- = 1 is the same line 68 68 there are many "forms" for the equation of a line to take; and all of them are equally valid

OpenStudy (amistre64):

....and I missed the -7 in my head lol